OpenStudy (anonymous):

Find the slope of each line. Through (-13,8) and (17, -6) My answer ended up being 2/4 am I correct?

OpenStudy (anonymous):

y2-y1 over x2-x1 plug it in and its that easy :)

OpenStudy (anonymous):

RISE/RUN

OpenStudy (pokemon23):

used this website http://www.wolframalpha.com

OpenStudy (pokemon23):

this helps you solve any problem

OpenStudy (pokemon23):

add that website yo your favorite

OpenStudy (anonymous):

How I did it was y2-y1(-6-8) over x2-x1(17- -13) and ended up with the answer 2/4 so do I flip it or am I good?

OpenStudy (anonymous):

you are good!

OpenStudy (anonymous):

-6-8=-14 17+13=30 -14/30=...